At what distance from the central axis of a long straight thin wire carrying a current of 5.0 A is the magnitude of the magnetic field due to the wire equal to the strength of the Earth's magnetic field of about 5.0 × 10-5 T? (μ0 = 4π × 10-7 T ∙ m/A)


1
Expert's answer
2021-07-27T07:02:01-0400

"B =\frac{ μ₀I}{2πr}"

"B = 5.0 \u00d7 10^{-5} T"

"I = 5.0A"

"\u03bc\u2080 = 4\u03c0\u00d710^{-7}T.m\/A"

"5.0 × 10^{-5}= \frac{4π × 10^{-7} T ∙ m/A × 5.0}{2πr}"

"5.0 \u00d7 10^{-5} \u00d7 2\u03c0r = {4\u03c0\u00d710 ^\n{\u22127}\n T\u2219m\/A\u00d75.0}"

"r = \frac{4π × 10^{-7}m}{2π×10^{-5}}"

"r = 2cm"/ "0.02m"